# Towards Code Generation for Octree-Based Multigrid Solvers

### v1

@@ 2 Software
To generate multigrid solvers on domains with mesh refinement, data structures like octrees had to be made accessible to the code generation pipeline. These data structures are usually highly complex and are often implemented in software frameworks written in general-purpose languages (GPLs) such as C/C++. We bridge the gap by acquainting the code generator with mesh refinement concepts from other frameworks, making them accessible for the generated solvers via automated interfaces. This approach simplifies the generation process by utilizing existing implementations for complex data structures. This work builds upon an existing coupling strategy between the waLBerla and ExaStencils frameworks [1].
ExaStencils [2] offers the generation of whole programs in C++ for geometric multigrid solvers from its own external domain-specific language (DSL) called ExaSlang. ExaSlang follows a multi-layered language approach, with each of the four layers tailored for domain experts from different communities [3]. ExaStencils employs a Scala-based source-to-source compiler to parse ExaSlang sources, apply code transformations for automatic parallelization and optimization, and pretty-print C++ code. The generated code supports automatic parallelization with OpenMP, MPI, and CUDA, along with optimizations such as common subexpression elimination (CSE), address pre-calculation, SIMD vectorization, and polyhedral loop transformations.
waLBerla [4] is a modular C++17 framework dedicated to computational fluid dynamics (CFD), especially for the Lattice Boltzmann method (LBM). Code generation of single kernels is enabled via the Python modules pystencils and lbmpy. pystencils, leveraging SymPy, generates optimized stencil codes with automatic SIMD vectorization, CSE, and non-temporal stores. lbmpy extends pystencils with further abstractions for the generation of efficient LBM kernels. waLBerla supports MPI for distributed-memory parallelism and OpenMP for shared-memory parallelism. The execution on AMD and NVIDIA GPU platforms is facilitated via HIP and CUDA [5].
@@
ExaStencils employs a hierarchical domain partitioning approach inspired by parallelism across hardware layers and supports a spectrum of grids ranging from uniform to non-uniform grids [6, 7]. Domains are organized into blocks and fragments. Blocks, the top hierarchy level, are mapped to distributed-memory parallel entities like MPI processes. Each block can be divided into equally sized fragments, each exclusively linked to a block. Fragments represent parts of the computational grid and are often mapped to sockets, cores, or accelerators, facilitating another layer of parallelization.
In waLBerla, an octree-based domain partitioning is used [4]. Initially, a cuboidal domain is regularly divided into equally-sized sub-domains, each serving as the root of an octree. These root blocks collectively form the blockforest, waLBerla’s central data structure, and can undergo further recursive refinement. Figure 1 illustrates this procedure on a two-dimensional unit domain $\Omega$ with four root blocks being incrementally refined. The refinement levels are depicted by colored contours, with red representing level three, green level two, and blue level one. Starting from a refinement level of zero on the left-most side, all root blocks are first globally refined once, followed by a local refinement in $\Omega_{1}$ and $\Omega_{2}$, respectively. The refinement levels are depicted by colored contours, with red representing level three, green level two, and blue level one. These refinement selection steps are commonly specified in user-defined functions or in custom configuration files. Once the refinement is completed, the blocks at the leaves of this structure represent the final sub-domains. In the context of distributed memory parallelism, each block is assigned exclusively to one process, although a process may manage multiple blocks. The only restriction for the refinement is a 2:1 size ratio between adjacent blocks.
@@ 3.1 Data Exchange
In this work, we will put high emphasis on the C2F and F2C cases. The same-level ghost layer exchange from [6] remains unchanged. From Fig. 2, it becomes evident that coinciding cell regions between the coarse block and its adjacent fine blocks must be determined for both cases. For a block with level $L_{curr}$ and neighbor in one of the cardinal directions with level $L_{neigh}$, a refinement ratio $r$ and dimensionality $D$, we define the number of potential neighbors as
@@ 4.1 Data Structures
Introducing mesh refinement concepts to an existing codebase causes major changes to the core data structures. Without mesh refinement, the number of potential neighbor blocks per dimension is known beforehand which allows for a static allocation of data structures for geometric and topologic information. In ExaStencils, this information is stored in so-called internal variables (IVs) [7]. One example for them is a boolean variable storing whether an adjacent fragment is on a remote MPI process and is used to distinguish between remote and local communication schemes. To support refined meshes with a variable number of blocks per process and a variable number of neighbors, we now rely on dynamic data structures such as C++ standard vectors. Topological information is now stored in an additional array dimension for the $N_{neigh}$ neighbor blocks. Within our generator, we introduced wrapper data structures with dynamic data allocation that are automatically employed in the case of mesh refinement. In our software coupling, these data structures are set up when the interface between the generated solver and the framework code is initialized.
@@ 4.2 Operators
Operators for numerical solvers usually require geometric information such as the cell width of a block as shown in the operator declaration for a gradient approximation in LABEL:lstlisting:grad_decl. Depending on the properties of a block, this information can either be expressed with a simple expression (or even a constant value) in the case of uniform blocks or must be stored explicitly in the case of non-uniform blocks. For this purpose, ExaStencils provides abstractions in the DSL, the so-called virtual fields [8], that obfuscate storage details and let the generator choose the most efficient case to obtain this information. Despite using uniform building blocks, having locally refined resolutions implies having a non-uniform discretization of the domain, which requires the operators to also consider refinement information. In this work, we want to go beyond the existing concepts and establish an automatic re-discretization of an operator for refinement without changing the user interface. We achieved that by implementing wrapper data structures for virtual fields that automatically fetch this information via accessors provided by a waLBerla block.
@@ 5 Application
In this chapter, we want to investigate the implications of mesh refinement on the numerical accuracy of multigrid solvers. Preliminary experiments have shown that the order of the C2F interpolation plays a vital role in error convergence. We showcase our proposed C2F interpolation scheme for a simple Poisson problem and compare it with a reference C2F scheme from waLBerla that employs a constant-order interpolation.
Consider the PDE for the three-dimensional Poisson problem given by
$-\Delta u(x,y,z)=f(x,y,z),\quad x,y,z\text{ on }\Omega$
where $\Delta$ is the Laplacian operator and $f(x,y,z)=0$ is a given right-hand side function. The domain is defined as $\Omega=[0,1]^{3}$ and Dirichlet boundary conditions have been set as
$u(x,y,z)=x^{2}-0.5y^{2}-0.5z^{2},\quad x,y,z\text{ on }\partial\Omega.$
We then discretize the continuous problem using the cell-centered finite volume method (FVM) and solve the arising linear system of equations with a $V(3,3)$ multigrid solver. The solver employs a damped Jacobi smoother ($\omega=0.8$) that is also used as the coarse-grid solver. After each multigrid cycle, a check for a $L_{2}$ residual norm threshold of $1\mathrm{E}{-16}$ or a maximum of 25 iterations is performed. Once converged, we obtain an approximate solution $u^{*}(x,y,z)$ that is expected to match the analytical formula from the boundary conditions. For this study, we compare the $L_{2}$ error norm results computed with $\left\|u^{*}-u\right\|_{2}$ for different resolutions. This is done by employing the same block partitioning of the domain as shown in Fig. 7 with 56 blocks on refinement level 2 (shaded in blue) and 64 blocks on refinement level 3 (shaded in red). Their resolution is then increased by raising the number of cells per block. The number of cells listed is regarded as the finest multigrid level which is then coarsened twice to form coarser multigrid levels. We employ a second-order discretization of the PDE and hence expect the error reduction rate
$\kappa=\frac{\|u^{*}_{h/2}-u_{h/2}\|_{2}}{\|u^{*}_{h}-u_{h}\|_{2}}$
to be $0.25$ when halving the step size $h$.
Figure 7: Domain with local refinement in center.
@@
$16^{3}$ 0.0105 -- 0.0001 -- 1.812e-05 --
$32^{3}$ 0.0103 0.9815 4.976e-05 0.476 4.543e-06 0.2507
$64^{3}$ 0.0102 0.9905 2.433e-05 0.489 1.137e-06 0.2502
$128^{3}$ 0.0101 0.9952 1.2036e-05 0.495 2.842e-07 0.2500
Table 1: $L_{2}$ error norm experiments over multiple resolutions with different interpolation orders.
As evident in Tab. 1, the C2F interpolation order strongly affects the error convergence of the numerical solver. The interpolation error of the constant-order C2F reference scheme appears to be the dominating factor for convergence since the error accuracy roughly remains the same for different resolutions. With our flexible approach, we generate a first- and second-order C2F scheme to further analyze the impact of the order on the convergence. The error reduction rate of the first-order scheme shows accuracy improvements for increasing the resolution but is limited to a linear factor. Our proposed second-order scheme, on the other hand, maintains the expected error reduction rate of the solver
@@ 8 Future Work
Another avenue is extending the mesh refinement implementation in the code generator with new features. This includes the support for other variable localizations than cell-centering. Optimizing the packing of our C2F schemes to employ a ring-buffer data structure for reusing intermediate extrapolation results for neighboring cells would decrease the time spent in communication even further. This can be done by writing code transformations to prepare the loop nests for the interpolation for the loop-carried CSE optimization from [9]. Achieving performance portability by supporting GPU platforms is another important milestone in the future and calls for platform-dependent optimizations such as using shared memory for storing reusable intermediate results.

### v2

@@ 2 Software
To generate multigrid solvers on domains with mesh refinement, data structures like octrees had to be made accessible to the code generation pipeline. These data structures are usually highly complex and are often implemented in software frameworks written in general-purpose languages (GPLs) such as C/C++. We bridge the gap by acquainting the code generator with mesh refinement concepts from other frameworks, making them accessible for the generated solvers via automated interfaces. This approach simplifies the generation process by utilizing existing implementations for complex data structures. This work builds upon an existing coupling strategy between the waLBerla and ExaStencils frameworks [exa_wb_coupling].
ExaStencils [exa_p2] offers the generation of whole programs in C++ for geometric multigrid solvers from its own external domain-specific language (DSL) called ExaSlang. ExaSlang follows a multi-layered language approach, with each of the four layers tailored for domain experts from different communities [exaslang]. ExaStencils employs a Scala-based source-to-source compiler to parse ExaSlang sources, apply code transformations for automatic parallelization and optimization, and pretty-print C++ code. The generated code supports automatic parallelization with OpenMP, MPI, and CUDA, along with optimizations such as common subexpression elimination (CSE), address pre-calculation, SIMD vectorization, and polyhedral loop transformations.
waLBerla [walberla] is a modular C++17 framework dedicated to computational fluid dynamics (CFD), especially for the Lattice Boltzmann method (LBM). Code generation of single kernels is enabled via the Python modules pystencils and lbmpy. pystencils, leveraging SymPy, generates optimized stencil codes with automatic SIMD vectorization, CSE, and non-temporal stores. lbmpy extends pystencils with further abstractions for the generation of efficient LBM kernels. waLBerla supports MPI for distributed-memory parallelism and OpenMP for shared-memory parallelism. The execution on AMD and NVIDIA GPU platforms is facilitated via HIP and CUDA [thermocapillary].
@@
ExaStencils employs a hierarchical domain partitioning approach inspired by parallelism across hardware layers and supports a spectrum of grids ranging from uniform to non-uniform grids [exa_automatic, Kuckuk2019]. Domains are organized into blocks and fragments. Blocks, the top hierarchy level, are mapped to distributed-memory parallel entities like MPI processes. Each block can be divided into equally sized fragments, each exclusively linked to a block. Fragments represent parts of the computational grid and are often mapped to sockets, cores, or accelerators, facilitating another layer of parallelization.
In waLBerla, an octree-based domain partitioning is used [walberla]. Initially, a cuboidal domain is regularly divided into equally-sized sub-domains, each serving as the root of an octree. These root blocks collectively form the blockforest, waLBerla’s central data structure, and can undergo further recursive refinement. Figure 1 illustrates this procedure on a two-dimensional unit domain $\Omega$ with four root blocks being incrementally refined. The refinement levels are depicted by colored contours, with red representing level three, green level two, and blue level one. Starting from a refinement level of zero on the left-most side, all root blocks are first globally refined once, followed by a local refinement in $\Omega_{1}$ and $\Omega_{2}$, respectively. The refinement levels are depicted by colored contours, with red representing level three, green level two, and blue level one. These refinement selection steps are commonly specified in user-defined functions or in custom configuration files. Once the refinement is completed, the blocks at the leaves of this structure represent the final sub-domains. In the context of distributed memory parallelism, each block is assigned exclusively to one process, although a process may manage multiple blocks. The only restriction for the refinement is a 2:1 size ratio between adjacent blocks.
@@ 3.1 Data Exchange
In this work, we will put high emphasis on the C2F and F2C cases. The same-level ghost layer exchange from [exa_automatic] remains unchanged. From Fig. 2, it becomes evident that coinciding cell regions between the coarse block and its adjacent fine blocks must be determined for both cases. For a block with level $L_{curr}$ and neighbor in one of the cardinal directions with level $L_{neigh}$, a refinement ratio $r$ and dimensionality $D$, we define the number of potential neighbors as
@@ 4.1 Data Structures
Introducing mesh refinement concepts to an existing codebase causes major changes to the core data structures. Without mesh refinement, the number of potential neighbor blocks per dimension is known beforehand which allows for a static allocation of data structures for geometric and topologic information. In ExaStencils, this information is stored in so-called internal variables (IVs) [Kuckuk2019]. One example for them is a boolean variable storing whether an adjacent fragment is on a remote MPI process and is used to distinguish between remote and local communication schemes. To support refined meshes with a variable number of blocks per process and a variable number of neighbors, we now rely on dynamic data structures such as C++ standard vectors. Topological information is now stored in an additional array dimension for the $N_{neigh}$ neighbor blocks. Within our generator, we introduced wrapper data structures with dynamic data allocation that are automatically employed in the case of mesh refinement. In our software coupling, these data structures are set up when the interface between the generated solver and the framework code is initialized.
@@ 4.2 Operators
Operators for numerical solvers usually require geometric information such as the cell width of a block as shown in the operator declaration for a gradient approximation in LABEL:lstlisting:grad_decl. Depending on the properties of a block, this information can either be expressed with a simple expression (or even a constant value) in the case of uniform blocks or must be stored explicitly in the case of non-uniform blocks. For this purpose, ExaStencils provides abstractions in the DSL, the so-called virtual fields [Kuckuk2017towards], that obfuscate storage details and let the generator choose the most efficient case to obtain this information. Despite using uniform building blocks, having locally refined resolutions implies having a non-uniform discretization of the domain, which requires the operators to also consider refinement information. In this work, we want to go beyond the existing concepts and establish an automatic re-discretization of an operator for refinement without changing the user interface. We achieved that by implementing wrapper data structures for virtual fields that automatically fetch this information via accessors provided by a waLBerla block.
@@ 5 Application
This chapter investigates the numerical accuracy of multigrid solvers on octrees with different C2F schemes. We consider a standard benchmark problem given by the 3D Poisson equation (Eq. 1) with the Laplacian operator $\Delta$, right-hand side function $f(x,y,z)=0$, and Dirichlet boundary conditions (Eq. 2).
$-\Delta u(x,y,z)=f(x,y,z)$ $\qquad x,y,z\text{ on }\Omega$ (1)
$u(x,y,z)=sin(\pi\cdot x)\cdot sin(\pi\cdot y)\cdot sinh(\sqrt{2}\cdot\pi\cdot z)$ $\qquad x,y,z\text{ on }\partial\Omega$ (2)
We discretize the continuous problem on the computational domain $\Omega=[0,1]^{3}$ using the cell-centered finite volume method (FVM) and solve the arising linear system with a $V(3,3)$ multigrid solver [Kuckuk2019]. The solver uses a damped Jacobi smoother with $\omega=0.8$, serving also as the coarse-grid solver. We terminate the solver when the $L_{2}$ residual norm falls below a threshold of $1\mathrm{E}{-16}$, or when the maximum of 35 iterations is reached. The obtained numerical solution $u^{*}$ is then used to compute the $L_{2}$ error norm as $\left\|u^{*}-u\right\|_{2}$. We compare errors across different resolutions using the block partitioning visualized in Fig. 7: it consists of 56 blocks on refinement level 2 (blue) and 64 blocks on level 3 (red).
Figure 8 summarizes the obtained results. The given block size is equal to the number of cells per block on the finest multigrid level. Coarser levels are added until a local resolution of $4^{3}$ cells is reached. Apart from the error norm itself, its grid convergence $\kappa$ (Eq. 3) is a vital metric. We expect quadratic convergence, that is when halving the local cell size $h$, $\kappa$ should be $0.25$, as we employ a second-order discretization of our PDE.
$\kappa=\frac{\|u^{*}_{h/2}-u_{h/2}\|_{2}}{\|u^{*}_{h}-u_{h}\|_{2}}$ (3)
We can see two effects. One, the absolute error decreased with increasing interpolation orders. Second, and more importantly, the chosen interpolation order limits the convergence order (constant for constant, linear for first-order and quadratic for second-order). To obtain the quadratic convergence expected from the PDE discretization, a second-order scheme is required. This further motivates the necessity to support higher-order interpolation schemes in communication routines although this usually entails additional implementation effort.
Figure 7: Domain with local refinement in the center.
@@
$16^{3}$ 0.235 -- 1.067e-2 -- 3.017e-3 --
$32^{3}$ 0.227 0.966 5.255e-3 0.493 7.215e-4 0.239
$64^{3}$ 0.224 0.984 2.611e-3 0.497 1.764e-4 0.244
$128^{3}$ 0.222 0.992 1.302e-3 0.499 4.361e-5 0.247
$256^{3}$ 0.221 0.996 6.499e-4 0.499 1.084e-5 0.249
Figure 8: $L_{2}$ error norm experiments over multiple resolutions with different interpolation orders.
@@ 8 Future Work
Another avenue is extending the mesh refinement implementation in the code generator with new features. This includes the support for other variable localizations than cell-centering. Optimizing the packing of our C2F schemes to employ a ring-buffer data structure for reusing intermediate extrapolation results for neighboring cells would decrease the time spent in communication even further. This can be done by writing code transformations to prepare the loop nests for the interpolation for the loop-carried CSE optimization from [Kronawitter2016]. Achieving performance portability by supporting GPU platforms is another important milestone in the future and calls for platform-dependent optimizations such as using shared memory for storing reusable intermediate results.